OpenStudy (anonymous):

integral from 2 to infinity of dx/x^2 - 1?

OpenStudy (kira_yamato):

Substitute u = sqrt(x) du/dx = 1 / (2sqrt(x)) dx = 2sqrt(x) du dx = 2u du Inverse function of u: x = sqrt(u) So: integral from 2 to infinity of dx/x^2 - 1 \[=\int\limits_{\sqrt{2}}^{\infty} \frac{2u}{u-1}du\] You can use integration by parts from here. Note: This is my opinion... Others may not agree